# The unified non-disciplinary model of the operation of neurons

### v3

@@ 4.5 Two segments
Now, we split the volume into two segments and compose them from ’thin physical layers’ (sheets) with different potentials (and concentrations). We consider the immediate environment of the neuron’s membrane as adjacent parallel $(x,y)$ plates and find the electrical field’s $z$ component of those plates in the points as shown in Fig. 2. In classic electricity, the charge remains on the surface of the conducting layer, and the distribution is homogeneous (although the charge is concentrated on the surface and the electrical field is zero inside the conducting layer). Hence, the electrical field contributions cancel each other. In electrolyte segments, (a tiny fraction of) the molecules decompose into the ionized state (dissociate), and the created ions interact with a bounding membrane using a not entirely understood mechanism [37]. The rest of the dipole molecules become polarized but do not dissociate, forming virtual charges that allow the electrolytes to contain electrical fields. Although not explicitly stated, we assume that electrical double layers are present near the membrane, and ions from the dispersion medium are adsorbed onto the particle’s surface, depending on the medium’s chemical properties.
@@ 4.5.1 Charge layers in segments
We consider that the segment is composed of electrically conducting discs (the ions are free to move on the surface), and the charged discs’ contribution to the electrical field at point $P$ (see Fig.  2) can be calculated as known from the theory of electricity. Due to the symmetry in the $z$ direction, in a homogeneous solution, the resulting electrical field in the plane perpendicular to the $z$ direction is zero, as we have contributions of equal magnitude with opposite signs.
@@ 4.5.3 Condenser in dielectric material
Given that the same number of charged ions must be present on the two sides of the membrane, their resultant surface density $\sigma_{A}$ must be the same. In the classic measurement by HH, the sum of the concentrations of $K^{+}$ and $Na^{+}$ in the cytoplasm and extracellular fluid (in mM) is 380 and 390, respectively, while the concentration of the $Cl^{-}$ are 450 and 460; practically all equal, as our analysis has derived. The deviations are well within the uncertainty of the measured values. Notice that the effect is purely electrostatic, resulting in an asymmetric ion distribution; in that balanced state, no permeability is needed. If the membrane is (slightly) permeable, ions will move across the membrane until equilibrium is reached. The resulting potential difference is directly proportional to the concentration difference.
@@ 4.5.4 Leakage current
HH did a complete measurement, and derived precise measured data. However, their measurement was not accurate because their model was wrong. In a balanced state, no current flows, so no voltage could be generated; see Fig. 3 in [43]: the leakage current is insignificant. More precisely: low intensity slow current flows between nearby ion channels inside the membrane, without involving the AIS. For this reason, they assumed that a permanent leakage current $I_{L}$ was flowing on the membrane, and on its resistance, it generated a 42.5mV42.5\ mV voltage [25]. In HH’s picture, the question what ion constitutes the current, remains open. It was claimed that the only permeability pathways open at rest are $g_{leak}$ and $g_{K}$. It follows that at the resting membrane potential, the leakage current equals the potassium current. That is, at rest, two currents flow (without a driving force in a balanced state), one consisting of $K^{+}$ ions and another of one $leakage^{+}$ ions, one generates voltage while the other not, and neither of them changes the concentration either on the departure or on the arrival segment, neither the concentration of the other ions; defying also Nernst’s law. BTW: [44] discusses the operation of the membrane, in some modified HH approach: it replaces "leakage" current with "pump" currents. In the wrong model, such a modification does not matter.
The idea of leakage current was wrong: the voltage HH measured [25] is correct, but – as our model explains – it was generated by charge separation rather than a current through a distributed resistor. It is permanently present while ionic segments are on the two sides of the membrane, even when no current flows. Measuring the energy consumption of neuronal operation [45] confirms that no resting current, in the sense as HH used the concept, exists. The existence of leakage current would be against, among others, the overall metabolic efficiency of neuronal operation. The idea of leakage current, alone, undermines the credibility of the HH model as it contradicts the observed heat absorption [22] during AP.
@@ 4.6 Bridging the micro- and macro views
due to the dielectrical properties of the segment. This contribution arises from the dielectricity in the segment (for simplicity, we have neglected the relative permittivity). It is to be added to the value of the classic contribution, the field generated by the conducting plates, so the final multiplier is $4$. (The experience is known in technical electricity: the electrolytic capacitors achieve several times higher charge storage capacity by using (pseudocapacitance). Using "roughened anode foil" increases the electrolyte thickness, and the roughening provides the necessary mechanical support. A thicker electrolyte layer wraps the condenser plates, and the dipoles in the thicker electrolyte provide an additional charge storage facility.)
@@ 4.10.4 Ion selectivity
The mechanism of channel passing can also contribute to explaining ion selectivity. "The normal selectivity cannot be explained by pore size, because $Na^{+}$ is smaller than $K^{+}$ [41]". The fact that ion selectivity changes with the membrane potential is observed (under the name of changed specific permeability), but it remains unexplained. "For example, the negative resting potential of nerve cells is largely determined by a class of $K^{+}$ channels that are 100-fold more permeable to $K^{+}$ than to $Na^{+}$. In contrast, during the action potential a class of $Na^{+}$ channels is activated that are 10- to 20-fold more permeable to $Na^{+}$ than to $K^{+}$. Thus, a key to the great versatility of neuronal signaling is the regulated activation of different classes of ion channels, each of which is selective for specific ions." [17], page 101. However, the question of why the permeability changes and how the channels are activated remains open. It is remarkable that an ion-specific channel is permeable also for other ions, only different transmission intensity is observed.
@@
The commonly used picture of the operation of selectivity filters is surely wrong. The assumed mechanical operation of the pores is too slow: the assumed structural change needs $10^{-8}\ s$ and the ions passage time is about $10^{-11}\ s$ (furthermore, it must be repeated about $10^{3}$ times per passage). If a wrong ion is caught, it must be transported back to its departure side, through the right ions (against their repulsion), and the right ions must retry. Neither forward nor backward movement has an appropriate driving force in that picture. Furthermore, the transport capacity is not sufficient for the operation.
@@
In our model, as Eq. (6) states, the electrolyte separated by a permeable membrane is balanced when the electrical and thermodynamic forces are equal. The $E_{thermal}^{C_{k}}(d)$ "electrical field" depends on the chemical quality of the ion; the electrical field does not. That means the resulting driving forces acting on the different ions differ. While a resulting force is exerted on one ion and moves it to the other segment, the other remains resting, or even moves in the opposite direction. Furthermore, the rush-in of ions greatly increases the $Na^{+}$ concentration in the dynamic layer, naturally increasing the number of $Na^{+}$ ions passing through the channel. That increase alone creates the illusion of the increased $Na^{+}$ permeability. Of course, the ions moving across the segments change the field contributions, indirectly affecting the balanced state of the other ions. The permanently open ion channels participate passively in the process, and their apparent permeability changes with the applied fields. To provide channel penetrability data is meaningful only if ion concentrations and membrane thickness data are also available. The driving forces of the ions differ, and they produce, depending on the actual concentrations and local potential, the selectivity of the channels.
@@ 4.10.5 Ion pumps
It is not clear what a force can "transport ions against their electrical and chemical gradients". It is undoubtedly "not on the ground that there is any ‘new force’ or whatnot" [7] present: the ions, independently of whether they are in an inanimate or living environment, can be moved only by electrical or thermodynamic forces, instead of some magic protein mechanism (ion pumps). The resultant potential (the sum of the electrical and thermodynamic potential for the individual ions) moves the ions through the channels in the membrane. In a typical mammalian cell, see Table 1, the sum of the electrical and thermodynamic potentials is $32\ [mV]\ K^{+}$ ions and $4\ [mV]\ Na^{+}$ ions. The same values for squids are $33\ [mV]\ K^{+}$ and $13\ [mV]\ Na^{+}$. Correspondingly, in the resting state, the resulting force moves sodium ions out of the cell and potassium ions into the cell (for more numerical examples see section 6.1); the same way as the "Na-K pump" is supposed to do by using some hypothesized protein mechanisms. In the classical model, no mechanism is provided for how the ion pumps gain energy from ATP and convert it into forward force. In our model, the potential gradient near the membrane, the ATP, by hydrolysis, generates ions, which move to the "plates" of the condenser, thus increasing its potential energy. That energy is prepared "globally" (not at the proximity where it is used) and continuously (during the full life cycle of neuronal operation), mainly in the recovery period. As discussed in section 5.2, in the transient state, to restore the condenser’s voltage, those ions must be collected.
@@
In our model, the ion pumps are ion channels which work in stationary rather than transient conditions. In the resting state, since the resultant forces exerted on the different ions differ, they permanently pass through the membrane. As described, the passage changes the ionic gradients on the two sides of the membrane and generates a slight "leakage current" and slightly fluctuating voltage. In the resting state, the electrolyte layers have sufficient time for after-diffusion, allowing a graded ionic gradient to form near the membrane. (The gradient changes in a fraction of a nanometer, so it is challenging to measure it.) With reference to Fig. 7, the graded gradient generates an electrical field on the ion channels, which is by orders of magnitude lower than before rush-in. Conversely, in the transient mode, the capped ion channels enable the creation of ungraded (step-like) gradients. Consequently, the ion speed, furthermore, the charge and mass transport, are by order of magnitude lower in the resting state than in the transient state, as observed.
@@
As Figure 6 in [43] displays, the ratio of $K+$ and $Na+$ changes sharply during the transient state. Theoretically, they assumed that the neuron “pumps 3 $Na^{+}$ ions out of the cell and two potassium ions in”; experimentally, they show in their Fig. 6 that the ratio changes between 0.01 and 7.5. According to our theory, the magnitude of the current (and so the exchange ratio) depends on the resulting potentials, so the ratio must change as the membrane potential changes during AP. The classical theory cannot explain this behavior and leads to exciting conclusions: "Furthermore, we analyzed energy properties of each ion channel and found that, under the two circumstances, power synchronization of ion channels and energy utilization ratio have significant differences. This is particularly true of the energy utilization ratio, which can rise to above 100% during subthreshold activity." [43] Our model says that (due to $Na^{+}$ rush-in) both the concentration and local membrane potential drastically change during the transient state. As long as the resulting potential is above $\approx 33\ mV$, the driving force acting on potassium ions is positive, i.e., the pump will move both ions out of the cell.
@@ 4.10.7 Voltage sensing
In our model, a voltage-controlled ion channel gets opened and closed due to purely electrostatic reasons. It works as a two-plate simple nano-scale electrometer (of type quadrant, Lindermann, Hoffman, and Wulf), similar to the ones used to measure the small electrical potential between charged elements (e.g., plates or fine quartz fibers). Given that the membrane and the cap in their resting state are isolators, no electrical repulsion evokes between them, and the adhesion sticks them firmly together, representing a permanent force. The van der Waals force is inversely proportional to the squared distance between the dipoles in the cap and the membrane, respectively, and is linearly proportional to the perimeter of the channel.
@@ 4.10.8 Gated ion channels in action
The gate remains open as long as the local charge distribution enables it. The cap is connected to the membrane only at one point, so it cannot fly away, and also cannot close again as long as the charge on the surface is present. In the absence of charge, the cap undergoes random motion, and the short-range van der Waals force may eventually reposition it on the membrane, thereby closing the channel. The voltage-sensing electrometer opens the channel, and the lack of charge on the surface enables it to close, but the closure is not immediate (the cap’s mass is by orders of magnitude larger than the mass of an ion that can pass through the channel). The fluctuation of the voltage gradient due to the slow current in the layer in the proximity of the membrane near the ion channel’s exit opens, closes, and re-opens the channel in an apparently stochastic way (actually, as the repulsion of charges due to the fluctuating current on the cap and the membrane regulates it), as observed.

### v4

@@ 4.5 Two segments
Now, we split the volume into two segments and compose them from ’thin physical layers’ (sheets) with different potentials (and concentrations). We consider the immediate environment of the neuron’s membrane as adjacent parallel $(x,y)$ plates and find the electrical field’s $z$ component of those plates in the points as shown in Fig. 3. In classic electricity, the charge remains on the surface of the conducting layer, and the distribution is homogeneous (although the charge is concentrated on the surface and the electrical field is zero inside the conducting layer). Hence, the electrical field contributions cancel each other. In electrolyte segments, (a tiny fraction of) the molecules decompose into an ionized state (dissociate), and the created ions interact with a bounding membrane using a not entirely understood mechanism [54]. The rest of the dipole molecules become polarized but do not dissociate, forming virtual charges that allow the electrolytes to contain electrical fields. Although not explicitly stated, we assume that electrical double layers are present near the membrane, and ions from the dispersion medium are adsorbed onto the particle’s surface, depending on the medium’s chemical properties.
@@ 4.5.1 Charge layers in segments
We consider that the segment is composed of electrically conducting discs (the ions are free to move on the surface), and the charged discs’ contribution to the electrical field at point $P$ (see Fig.  3) can be calculated as known from the theory of electricity. Due to the symmetry in the $z$ direction, in a homogeneous solution, the resulting electrical field in the plane perpendicular to the $z$ direction is zero, as we have contributions of equal magnitude with opposite signs.
@@ 4.5.2 Simple conducting sheet (a physical condenser)
and at a 5nm5\ nm membrane thickness, the voltage on a classical (non-dielectric) condenser across the plates becomes
@@ 4.5.3 Condenser in dielectric material
Given that the same number of charged ions must be present on the two sides of the membrane, their resultant surface density $\sigma_{A}$ must be the same. In the classic measurement by HH, the sum of the concentrations of $K^{+}$ and $Na^{+}$ in the cytoplasm and extracellular fluid (in mM) is 380 and 390, respectively, while the concentration of the $Cl^{-}$ are 450 and 460; practically all equal, as our analysis has derived they should be. The deviations are well within the uncertainty of the measured values. Notice that the effect is purely electrostatic, resulting in an asymmetric ion distribution; in that balanced state, no permeability is needed. If the membrane is (slightly) permeable, ions will move across the membrane until equilibrium is reached. The resulting potential difference is directly proportional to the concentration difference.
@@ 4.5.4 Leakage current
HH did a complete measurement, and derived precise measured data. However, their measurement was not accurate because their model was wrong. In a balanced state, no current flows, so no voltage could be generated; see Fig. 3 in [36]: the leakage current is insignificant. More precisely: low intensity slow current flows between nearby ion channels inside the membrane, without involving the AIS. For this reason, they assumed that a permanent leakage current $I_{L}$ was flowing on the membrane, and on its resistance, it generated a 42.5mV42.5\ mV voltage [37]. In HH’s picture, the question what ion constitutes the current, remains open. It was claimed that the only permeability pathways open at rest are $g_{leak}$ and $g_{K}$. It follows that at the resting membrane potential, the leakage current equals the potassium current. That is, at rest, two currents flow (without a driving force in a balanced state), one consisting of $K^{+}$ ions and another of $leakage^{+}$ ions, and one generates voltage.At the same time, neither of them changes the concentration on either the departure or the arrival segment, nor the concentration of the other ions, defying also Nernst’s law. BTW: [58] discusses the operation of the membrane, in some modified HH approach: it replaces "leakage" current with "pump" currents. In the wrong model, such a modification does not matter.
The idea of leakage current was wrong: the voltage HH measured [37] is correct, but – as our model explains – it was generated by charge separation rather than a current through a distributed resistor. It is permanently present while ionic segments are on the two sides of the membrane, even when no current flows. Measuring the energy consumption of neuronal operation [35] confirmed that no resting current, in the sense as HH used the concept, exists. The existence of leakage current would be against, among others, the overall metabolic efficiency of neuronal operation.
The idea of leakage current, alone, undermines the credibility of the HH model as it contradicts the observed heat absorption [26] during AP. The lack of leakage current not only validates our model, but also solves the long-standing mystery of reversible heat release during the AP.
The energy consumption of the nervous impulse underpins that the complex physical process comprises mostly reversible disciplinary processes: most of the energy of AP is stored as reversible elastic potential energy. Hence, its dissipation is also negligible: "no significant net heat is produced" [40].
Alternatively, it may be hypothesized that the positive $K^{+}$ ions, while traveling ahead of the positive $Na^{+}$ ions under the effect of a potential-free space in the second phase of AP, not only turn with their positive charge the offset voltage to negative, but on the way the $K^{+}$ ions collect the heat dissipated by the $Na^{+}$ ions. "Laws of science cannot describe life". Heat liberation and absorption are direct experimental evidence of the "slow current" and the mechanisms described above.
The shocking conclusion in [59], that "brain cells communicate with mechanical pulses, not electrical signals", more precisely sounds that mechanical pulses (pressure waves) deliver the charge carriers of the electrical signals, in close interaction with the repulsion of the delivered charged carriers. In Equation (9), we also must consider the mechanical force delivered by the pressure wave. Furthermore, the nervous pulse transmission over the axons must be revisited. The classical mechanism based on the coordinated action of ion channels is undoubtedly wrong. The enormous pressure, the slightly compressible electrical fluid, and the repulsion of the delivered particles suggest an alternative and more reasonable mechanism.
@@ 4.6 Bridging the micro- and macro views
due to the dielectric properties of the segment. This contribution arises from the dielectricity in the segment (for simplicity, we have neglected the relative permittivity). It is to be added to the value of the classic contribution, the field generated by the conducting plates, so the final multiplier is $4$. (The experience is known in technical electricity: the electrolytic capacitors achieve several times higher charge storage capacity by using pseudocapacitance. Using "roughened anode foil" increases the electrolyte thickness, and the roughening provides the necessary mechanical support. A thicker electrolyte layer wraps the condenser plates, and the dipoles in the thicker electrolyte provide an additional charge storage facility.)
@@ 4.10.4 Ion selectivity
"Some types of cation-selective channels allow the cations that are usually present in extracellular fluid – $Na^{+}$, $K^{+}$, $Ca^{2+}$, and $Mg^{2+}$ – to pass almost indiscriminately. However, many other cation- selective channels are permeable primarily to a single type of ion, whether it is $Na^{+}$, $K^{+}$, or $Ca^{2+}$." "The normal selectivity cannot be explained by pore size, because $Na^{+}$ is smaller than $K^{+}$[27], page 107. The fact that ion selectivity changes with the membrane potential is observed (under the name of changed specific permeability), but it remains unexplained. "For example, the negative resting potential of nerve cells is largely determined by a class of $K^{+}$ channels that are 100-fold more permeable to $K^{+}$ than to $Na^{+}$. In contrast, during the action potential a class of $Na^{+}$ channels is activated that are 10- to 20-fold more permeable to $Na^{+}$ than to $K^{+}$. Thus, a key to the great versatility of neuronal signaling is the regulated activation of different classes of ion channels, each of which is selective for specific ions." [19], page 101. However, the question of why the permeability changes and how the channels are activated remains open. It is remarkable that an ion-specific channel is permeable also for other ions, only different transmission intensity is observed. Unfortunately, the measurements are incomplete: neither the time (say, relative to the last AP), nor the neuron’s state, nor the local potentials at the two ends of the channel (at the two dynamic layers), nor the location of the channel on the membrane are provided.
@@
The above mechanism of channel passage can also help explain ion selectivity. The commonly used picture of the operation of selectivity filters is surely wrong. The assumed mechanical operation of the pores is too slow: the assumed structural change needs $10^{-8}\ s$ and the ions passage time is about $10^{-11}\ s$ (furthermore, it must be repeated about $10^{3}$ times per open period). If a wrong ion is caught, it must be transported back to its departure side, through the right ions (against their repulsion), and the right ions must retry. Neither forward nor backward movement has an appropriate driving force in that picture (a ’whatnot’ force). Furthermore, the transport capacity is not sufficient for that kind of operation.
@@
In our model, as Eq. (8) states, the electrolyte separated by a permeable membrane is balanced when the electrical and thermodynamic forces are equal. The $E_{thermal}^{C_{k}}(d)$ "electrical field" depends on the chemical quality of the ion; the electrical field does not. That means the resulting driving forces acting on the different ions differ. While a resulting force is exerted on one ion and moves it to the other segment, the other remains resting, or even moves in the opposite direction. Furthermore, the rush-in of ions greatly increases the $Na^{+}$ concentration in the dynamic layer, naturally increasing the number of $Na^{+}$ ions passing through the channel. That increase alone creates the illusion of the increased $Na^{+}$ permeability. Of course, the ions moving across the segments change the field contributions, indirectly affecting the balanced state of the other ions. The permanently open ion channels participate passively in the process, and their apparent permeability changes with the applied fields. To provide channel penetrability data is meaningful only if ion concentrations and membrane thickness data are also available. The driving forces of the ions differ, and they produce, depending on the actual concentrations and local potential, the selectivity of the channels. Of course, the driving force is a statistical one: it is not valid for the individual ions, but it is valid for the result of a population of ion channels.
@@ 4.10.5 Ion pumps
It is not clear what a force can "transport ions against their electrical and chemical gradients". It is undoubtedly "not on the ground that there is any ‘new force’ or whatnot" [9] present: the ions, independently of whether they are in an inanimate or living environment, can be moved only by electrical or thermodynamic forces, instead of some magic protein mechanism (ion pumps). The resultant potential (the sum of the electrical and thermodynamic potential for the individual ions) moves the ions through the channels in the membrane. In a typical mammalian cell, in the resting state, see Table 1, the sum of the electrical and thermodynamic potentials is $32\ [mV]\ K^{+}$ ions and $4\ [mV]\ Na^{+}$ ions. The same values for squids are $33\ [mV]\ K^{+}$ and $13\ [mV]\ Na^{+}$. Correspondingly, in the resting state, the resulting force moves sodium ions out of the cell and potassium ions into the cell (for more numerical examples see section 6.1). The resultant potential difference actually implements what "Na-K pump" is supposed to do by using some hypothesized protein mechanisms.
In the classical model, no mechanism is provided for how the ion pumps gain energy from ATP and convert it into forward force. In our model, the potential gradient near the membrane, the ATP, by hydrolysis, generates ions, which move to the "plates" of the condenser, thus increasing its potential energy. Claims such as "The flux of ions through ion channels is passive, requiring no expenditure of metabolic energy by the channels." [19], page 107, are wrong. The enormous electric field across the membrane performs work on the ion, and the effect of the transmitted ion decreases the accelerating potential: some (white) ions are released, and those are replaced by ions produced by the hydrolysis of ATP as discussed throughout the paper. There is no "free passive transport". That energy is prepared "globally" (not at the proximity where it is used) and continuously (during the full life cycle of neuronal operation), mainly in the recovery period. As discussed in section 5.2, in the transient state, to restore the condenser’s voltage, those ions must be collected.
@@
In our model, the ion pumps are ion channels which work in stationary rather than transient conditions. In the resting state, the potential difference fluctuates around zero (due to the permanent slow material transport). In contrast, while at the beginning of the transient state, a sudden change around the entrance of the channel creates a large potential difference, so an intense transport happens. In the resting state, since the resultant forces exerted on the different ions differ, they permanently pass through the membrane. As described, the passage changes the ionic gradients on the two sides of the membrane and generates a slight "leakage current" and slightly fluctuating voltage. In the resting state, the electrolyte layers have sufficient time for after-diffusion, allowing a graded ionic gradient to form near the membrane. (The gradient changes in a fraction of a nanometer, so it is challenging to measure it.) With reference to Fig. 8, the graded gradient generates an electrical field on the ion channels, which is by orders of magnitude lower than before rush-in. Conversely, in the transient mode, the capped ion channels enable the creation of ungraded (step-like) gradients. Consequently, the ion speed, furthermore, the charge and mass transport, are by two orders of magnitude lower in the resting state than in the transient state, as observed, mainly because the conductance of the ion channels used in the two states differs by two orders of magnitude.
@@
As Figure 6 of [36] displays, the ratio of $K+$ and $Na+$ changes sharply during the transient state. Theoretically, they assumed that the neuron “pumps 3 $Na^{+}$ ions out of the cell and two potassium ions in”; experimentally, they showed in their Fig. 6 that the ratio changes between 0.01 and 7.5. According to our theory, the magnitude of the currents (and so the exchange ratio) depends on the resulting potentials, so the ratio must change as the membrane potential changes during AP. The classical theory cannot explain this behavior and leads to exciting conclusions: "Furthermore, we analyzed energy properties of each ion channel and found that, under the two circumstances, power synchronization of ion channels and energy utilization ratio have significant differences. This is particularly true of the energy utilization ratio, which can rise to above 100% during subthreshold activity." [36] (The simplest way to solve energy problems would be to use subthreshold excitations). Our model says that (due to $Na^{+}$ rush-in) both the concentration and local membrane potential drastically change during the transient state. As long as the resulting potential is above $\approx 33\ mV$, the driving force acting on potassium ions is positive, i.e., the pump will move both ions out of the cell. The composition of the output current depends on the input current.
@@ 4.10.7 Voltage sensing
In our model, a voltage-controlled ion channel gets opened and closed due to purely electrostatic reasons. It works as a two-plate simple nano-scale electrometer (of type quadrant ), similar to the ones used to measure the small electrical potential between charged elements (e.g., plates or fine quartz fibers). Given that the membrane and the cap in their resting state are isolators, no electrical repulsion evokes between them, and the adhesion sticks them firmly together, representing a permanent force. The van der Waals force is inversely proportional to the squared distance between the dipoles in the cap and the membrane, respectively, and is linearly proportional to the perimeter of the channel.
@@ 4.10.8 Gated ion channels in action
A gate is open when there is no mechanical obstacle against material transport, and the local resulting potentials on the two sides of the membrane in the near-membrane layers enable the transport of the ions. The cap is connected to the membrane only at one point, so it cannot fly away, and also cannot close again as long as the charge on the surface is present. In the absence of charge, the cap undergoes random motion, and the short-range van der Waals force may eventually reposition it on the membrane, thereby closing the channel. The voltage-sensing electrometer opens the channel, and the lack of charge on the surface enables it to close, but the closure is not immediate (the cap’s mass is by orders of magnitude larger than the mass of an ion that can pass through the channel). The fluctuation of the voltage gradient due to the slow current in the layer in the proximity of the membrane near the ion channel’s exit opens, closes, and re-opens the channel in an apparently stochastic way (actually, as the repulsion of charges due to the fluctuating current on the cap and the membrane regulates it), as observed.